# Image Generative Semantic Communication with Multi-Modal Similarity Estimation for Resource-Limited Networks

### v1

@@
To reduce network traffic and support environments with limited resources, a method for transmitting images with low amounts of transmission data is required. Several machine learning-based image compression methods, which compress the data size of images while maintaining their features, have been proposed. However, in certain situations, reconstructing the semantic information of images at the receiver end may be sufficient. To realize this concept, semantic-information-based communication, called semantic communication, has been proposed, along with an image transmission method using semantic communication. This method transmits only the semantic information of an image, and the receiver reconstructs the image using an image-generation model. This method utilizes one type of semantic information for image reconstruction, but reconstructing images similar to the original image using only it is challenging. This study proposes a multi-modal image transmission method that leverages diverse semantic information for efficient semantic communication. The proposed method extracts multi-modal semantic information from an original image and transmits only it to a receiver. Subsequently, the receiver generates multiple images using an image-generation model and selects an output image based on semantic similarity. The receiver must select the result based only on the received features; however, evaluating semantic similarity using conventional metrics is challenging. Therefore, this study explored new metrics to evaluate the similarity between semantic features of images and proposes two scoring procedures for evaluating semantic similarity between images based on multiple semantic features. The results indicate that the proposed procedures can compare semantic similarities, such as position and composition, between semantic features of the original and generated images. Thus, the proposed method can facilitate the transmission and utilization of photographs through mobile networks for various service applications, including monitoring, tracking, and detection.
@@ 1 Introduction
It is expected to increase situations wherein photos taken at specific locations are transmitted through a mobile network and used for various services and applications such as monitoring, tracking, and detection. Additionally, recent surveys of undersea resources using underwater communication [1, 2] and environmental monitoring using satellite communication in severe locations [3, 4, 5] have also been conducted. Therefore, image transmission technologies will have to realize transmitting numerous images from multiple locations in the future.
However, it is important to reduce network traffic owing to data transmission including image transmission, because global network traffic is expected to continue increasing [6, 7, 8]. Additionally, developing periodic image transmission techniques for recent applications in severe environments wherein communication resources are extremely limited such as underwater remains challenging. Consequently, image transmission using fewer network resources has the potential for use in various applications.
@@
These methods aim to transmit images while maintaining the visual identity of the images. However, transmitting all features of the images is redundant in certain situations. For example, in an application aimed toward monitoring specific objects, the background information in a captured photo is less important than that related to the target object, such as its type and composition. Moreover, when viewing a live camera stream at an event or exhibition, the faces of people and background information are not important for viewers because their primary interests lie in the event and exhibits. In such situations, the transmitter does not need to transmit the all features contained in the image to the receiver.
@@
Therefore, in such situations, if a transmitter can extract specific features from an original image based on the requirements of the receiver, the amount of data transmitted can be reduced by transmitting only the features extracted from the image. Additionally, owing to recent developments in the field of artificial intelligence, image-generation models [19, 20, 21] that allow generating images using descriptive information or sketches have become widely available. Moreover, recent real-time image-generation models [22] can apply image-generation mechanisms to time-sensitive applications. Consequently, there is potential for the development of image-generation-based image transmission, which can be considered a type of semantic communication in the future.
@@
As an example of such a kind of communications, the authors of [23] have proposed an image transmission method based on the descriptive information contained in an image. In this method, the transmitter transmits only the descriptive text information extracted from the original image using an image-to-text algorithm to the receiver, and the receiver then reconstructs the image based on the received information through an image-generation model. Although this method can significantly reduce the communication cost compared to existing image transmission methods, some important semantic features, such as object position and composition, might be lost because it only relies on a single semantic feature from the reconstructed image. Therefore, it is necessary to extract and transmit multi-modal semantic features from the original image to realize image reconstruction with high semantic similarity and satisfy the requirements of the receiver.
@@
To address this issue, we previously proposed an image transmission concept that employs multiple types of semantic information [24]. In this method, a transmitter extracts multiple semantic features from an original image, such as composition and color, in addition to captions, and transmits them to a receiver, which reconstructs the image using an image-generation model. The images generated using multiple semantic features are expected to have a higher similarity to the original images than those generated using only a single semantic feature. However, a quantitative analysis of the semantic similarity between the original and generated images was not conducted in [24].
@@
In this study, we propose a highly efficient image transmission method based on the concept proposed in [24] and employ newly designed evaluation metrics to assess the semantic similarity between the original and generated images. In the proposed method, an image-generation method at the receiver reconstructs images with the semantic information required by its application using the multi-modal semantic information extracted from the original image. Fig. 1 shows an overview of the proposed method. To achieve image transmission with a small amount of data, the transmitter transmits multiple features of the original image, and the receiver reconstructs an image with the same semantic information as the original image. First, the transmitter extracts multiple features representing composition, description, and color information from the original image as diverse semantic information using existing machine learning algorithms. Thereafter, the transmitter transmits only the extracted semantic features to the receiver, with the aim of minimizing the amount of data transmitted. The receiver generates multiple images by inputting the received features into an image-generation model. As the image-generation model generates various images, the receiver must select the best image as an output image. To evaluate the reconstruction result, this study developed quality assessment procedures, which is introduced later. In the proposed method, the receiver reconstructs images by the received semantic information to be easily recognized by humans. Therefore, the proposed method can be potentially used in various applications.
@@
To select an appropriate image using the proposed method, the quality of the generated images must be assessed based on the semantic similarity between the original and generated images on the receiver. Although some metrics assess the image quality by comparing two images [25], the proposed method cannot compare the original and generated images directly because the receiver does not possess the original image. Therefore, the quality of the generated images must be assessed based only on the received semantic features of the original image. To address this issue, this study investigates various metrics to evaluate the semantic similarity between the original and generated images, with a particular focus on multi-modal semantic information. This study proposes two scoring procedures for evaluating the similarity between semantic features, such as descriptive and segment information, of the original and generated images. Additionally, this study also proposes a background recoloring method for enhancing the contours of objects in images to improve image generation.
@@
  • We propose an image construction method that reconstructs images at the receiver end using multi-modal semantic communication. Therefore, the proposed method can transmit images with higher semantic similarity information than conventional transmission methods. Additionally, the proposed method only uses open-source machine learning algorithms, making it easy to deploy and underscoring its potential for use in various applications.
  • The proposed method demonstrates better reproducibility than [24] because it evaluates the generated images based on their similarity between the semantic features of the original and generated images. To realize this, we propose two scoring procedures for evaluating the semantic similarity between images. These procedures can evaluate the semantic similarity between images based only on the semantic features received by the receiver.
  • We propose and investigate new metrics for evaluating the semantic similarity between the original and generated images through an experiment. This is an important first step toward establishing evaluation metrics for image transmission methods using semantic communication. The results of the experiment indicate that the proposed scoring procedures can evaluate semantic features of the target object, such as position and composition, and the entire of background information of the image. Additionally, the results demonstrate that the proposed method with the background recoloring technique can effectively reconstruct the composition and position of the target object in the generated images.
@@ 2 Related Work
[23] proposed an image transmission method that transmits only the image caption to further improve communication efficiency. In this method, a transmitter generates a caption as descriptive information of the original image using an image captioning algorithm. Subsequently, the transmitter transmits only the generated caption, and the receiver reconstructs the image based on the received caption by using an image-generation model.
As this method extracts only the descriptive information from the original image, it may incur a loss of some semantic information contained in the original image that is difficult to represent through the descriptive information. For example, it is difficult to represent the exact position and composition of a target object in an image using only descriptive information. Therefore, to reconstruct an image more clearly at the receiver, the transmitter must extract and transmit multi-modal semantic information from the original image. To address this issue, we previously proposed a novel image transmission method that uses multi-modal semantic information [24]. In this method, a transmitter extracts descriptive, segmented, and color information from the original image as semantic features, and transmits only these features. After receiving these features, the receiver reconstructs the image using an image-generation model. By conveying diverse information, such as the composition and color of the original image, in addition to descriptive information, this method enables image transmission that can generate images that are more semantically similar to the original image compared with existing methods. However,  [24] included only subjective classification of the generated images as a preliminary study. Therefore, quantitative evaluations of the similarities between original and generated images are lacking.
@@ 3 Image Generation-based Transmission Method
This section describes the proposed image generation-based transmission method that uses multi-modal semantic information. A transmitter extracts multiple semantic features, such as composition, description, and color information, representing diverse semantic information from an image and transmits only these extracted features. The details of these extracted features are explained in Sec. 3.1. The receiver generates multiple images using the received features and selects an output image based on the semantic similarity. The details of these procedures at the receiver are explained in Sec. 3.2. The proposed method focuses on reducing both the amount of transmitted data and image reconstruction from the receiver’s perspective while maintaining the extracted semantic features.
@@ 4 Evaluation Metrics for the Proposed Method
Table 1: Data sizes employed by each method. The average data size of the captions is approximately $1/4000^{\textrm{th}}$ of that of the images in the JPEG format. Additionally, even when multiple semantic features are combined, the data size is approximately $1/20^{\textrm{th}}$ of that of the JPEG format images.
@@ 5 Experiment Setup
Refer to caption (a) Using only the caption. Refer to caption (b) Proposed method without background recoloring. Refer to caption (c) Proposed method with background recoloring.
@@
Refer to caption (a) Using only captions. Refer to caption (b) Proposed method without background recoloring. Refer to caption (c) Proposed method with background recoloring.
@@ 6.1 Data Sizes of Semantic Features
We compared the data size transmitted by the proposed method with that of the raw bit maps, JPEG formatted data, and generated captions of the original images to validate its communication efficiency. Table 1 presents the data sizes for each compression approach. “Uncompressed” denotes the size of the uncompressed bit map of the original images, “JPEG” denotes the average size of the original images in the JPEG format, and “Caption” denotes average size of the captions for each original image, which is used by the simple method employing only the descriptive information of the images. In addition, “Caption + color palette + segmentation array” denotes the average total size of the caption, color palette, and segmentation array for each image, which is the size of semantic features used in the proposed method. The sizes of all elements of the segmentation arrays and captions characters were calculated as 1 byte each. The data size of each color palette was calculated as the number of labels in an image $\times$ 4 bytes: 3 bytes for the RGB values and 1 byte for the label index. Each segmented array was compressed using run-length encoding for further data reduction. Additionally, “Caption + color palette + segmentation array” includes the RGB value of the “background” label.
The results indicate that the proposed methods achieved higher data reduction than the images in JPEG format, thereby underscoring the utility of the proposed method in situations with limited network resources. Additionally, data size comparisons of the semantic features showed that only the caption extracted from the images had a smaller data size than the total size of the multiple semantic features, indicating that the communication efficiency was better when only captions were used. However, the proposed method was superior in terms of color and composition reproduction, as described in the following sections.
@@ 6.3 Comparison of Segmentation Matching Rate of Each Method
Fig. 8 shows the segmentation matching rates for the segmentation arrays generated from the original 5 “airplane” images and the 50 images generated using each method. Approximately, the same trends can be observed in the results of the other images, wherein the proposed methods exhibit higher matching rates than using only captions for the image generation. This is because the proposed methods can generate images with similar compositions and positions of the target object as those in the original image using the colored-segmented images.
@@ 7 Conclusion
This paper proposed an image-generation-based transmission method. In the proposed method, a transmitter extracts multi-modal semantic features from an image and transmits them to a receiver, thereby significantly reducing the amount of transmitted data compared to traditional methods. After receiving the data, the receiver generates multiple images using an image-generation model and selects an output image based on the semantic similarity between the original and generated images. Therefore, the proposed method can realize both a significant reduction in the transmitted data size and reconstruction of the image with the semantic information required by the receiver. The evaluation results validated the ability of the proposed method to significantly reduce the data size compared with typical image compression algorithms.
In the proposed method, the receiver must compare and evaluate the similarity between the original and generated images using only the received data when selecting the output. Therefore, this study focused on descriptive and segment features of images and proposed two scoring methods for comparing them between the original and generated images.
@@
An experiment was conducted to verify the effectiveness of the proposed image transmission method and scoring procedures. The results indicated that the proposed method using multi-modal semantic information can generate images that are more similar to the original images than those using only a single piece of information. In particular, the proposed method exhibits good reconstruction performance for the composition and position of the objects in the original image because a colored-segmented image is input into the image-generation model as segment and color information. In addition, the proposed scoring procedures can quantify the semantic similarity between images. The text similarity scores can be used to compare the types of target objects and rough background information of images. In contrast, the segmentation matching rates can be used to compare the composition and position of the target object in the images. However, we found that if the target object in the image was too small, the generated images with no objects exhibited high segmentation matching rates.
@@
In a future work, we will improve the scoring procedures used for evaluating the semantic similarity between images at the receiver end. In the proposed scoring procedure for the segmentation matching rate, images generated with no objects had a high matching rate when the target object in the original image was too small. This is because the proposed procedure calculates the matching rate based on the number of pixels including those labeled as “background.” Therefore, it is necessary to develop a calculation method that considers only the pixels labeled as “target object.” Moreover, we plan to develop an image reconstruction method based on the evaluation results of semantic similarity between the images. In the proposed method, the receiver selects the output result from the generated images based on semantic similarity. However, if all generated images have low similarity scores, the receiver cannot select the result using the semantic information required by its application. To address this issue, it is necessary to develop a method to regenerate images or, if necessary, request the sender to transmit the semantic features again.

### v2

@@
To reduce network traffic and support environments with limited resources, a method for transmitting images with minimal transmission data is required. Several machine learning-based image compression methods, which compress the data size of images while maintaining their features, have been proposed. However, in certain situations, reconstructing only the semantic information of images at the receiver end may be sufficient. To realize this concept, semantic-information-based communication, called semantic communication, has been proposed, along with an image transmission method using semantic communication. This method transmits only the semantic information of an image, and the receiver reconstructs it using an image-generation model. This method utilizes a single type of semantic information for image reconstruction, but reconstructing images similar to the original image using only this information is challenging. This study proposes a multi-modal image transmission method that leverages various types of semantic information for efficient semantic communication. The proposed method extracts multi-modal semantic information from an original image and transmits only that to a receiver. Subsequently, the receiver generates multiple images using an image-generation model and selects an output image based on semantic similarity. The receiver must select the result based only on the received features; however, evaluating semantic similarity using conventional metrics is challenging. Therefore, this study explores new metrics to evaluate the similarity between semantic features of images and proposes two scoring procedures for evaluating semantic similarity between images based on multiple semantic features. The results indicate that the proposed procedures can compare semantic similarities, such as position and composition, between the semantic features of the original and generated images.
@@ 1 Introduction
1.1 Background
The application of photos taken in specific locations for services such as monitoring, detection, and tracking can be utilized in various scenarios. To realize such applications in multiple locations using lots of devices, sufficient communication resources are required within networks for periodically transmitting images. However, the following issues may cause a shortage of the communication resources within the networks: (1) Available communication resources are severely limited due to communication media with low data rates, and (2) the resource allocated to each device is limited because of utilizing communication media with high data rates by numerous devices.
The issue (1) may arise in severe environments where mobile networks are unavailable. Especially, methods for establishing networks in severe environments, such as underwater, mountain, and forest, attract attention in recent years for applications in monitoring, detecting and tracking  [1, 2, 3]. For example, underwater communication technologies using acoustic wave and visible light have been proposed for monitoring of underwater environments and surveys of undersea resources in underwater environments [1, 2]. For another example, low power wide area (LPWA) that realizes wide-area communication with low data rates attracts attention for monitoring and detecting in severe terrestrial environments such as mountains or forests [3, 4]. These communication technologies have a potential to realize the applications with image transmission in the severe environments. However, these technologies are difficult to periodically transmit large data such as images at multiple locations because their data rates are lower than those of mobile networks.
The issue (2) may arise in situations where numerous devices need to share abundant communication resources. Recently, global network traffic is expected to continue increasing due to an increase in the number of devices [5, 6, 7], and wireless communication technologies such as 5G and 6G rapidly advance to address this issue. In addition, new communication technologies such as satellite communication have been proposed to cover terrestrial areas without mobile networks and to realize non terrestrial network (NTN) [8, 4, 9]. However, the communication resources allocated to each device become exceedingly limited with an increase in the number of devices, even taking into account developments in these communication technologies. These trends may make image transmission in terrestrial and non terrestrial environments for monitoring, detecting, and tracking more difficult. Therefore, image transmission using fewer communication resources has the potential to be utilized in various applications across both urban and remote environments.
1.2 Semantic Communications
@@
These methods aim to compress and transmit images while maintaining the visual identity of the images. Therefore, these compressed images retain entire of the semantic information in the original images. However, in environments where communication resources are severely limited, as described in Sec. 1.1, it is difficult to transmit images compressed by these algorithms because the compressed images still have large data sizes. In addition, transmitting all features of the images is redundant in certain situations. For example, when monitoring or detecting vermin in natural environments such as large farms, it will be sufficient to reconstruct types and conditions at a receiver. In other words, in this situation, background information in images is less important than the information of the target objects. For another example, when monitoring or detecting specific objects such as vehicles using numerous devices that are allocated limited communication resources in urban environments, background information and facial information of people staying in the urban are not important. In these situations, it will be sufficient if only specific semantic information contained in the original images, such as presents, types, and compositions of the target objects, is reconstructed by the receiver.
@@
In such situations, if a transmitter can extract specific semantic features from an original image based on the requirements of the receiver, and the receiver can reconstruct the image with the same semantic features in the original one based on these features, the amount of the transmitted data be expected to significantly reduce by transmitting only the extracted features. Additionally, owing to recent developments in the field of artificial intelligence, image-generation models [20, 21, 22] that allow generating images using descriptive information or sketches have become widely available. Moreover, recent real-time image-generation models [23] can apply image-generation mechanisms to time-sensitive applications. Consequently, there is potential for the development of image-generation-based image transmission, which can be considered a type of semantic communication in the future.
@@
As an example of such a kind of communications, the authors of [24] have proposed an image transmission method based on the descriptive information contained in an image. In this method, the transmitter transmits only the descriptive text information extracted from the original image, and the receiver then reconstructs the image based on the received information through an image-generation model. Although this method can significantly reduce the communication cost compared to existing image transmission methods, some important semantic features, such as object position and composition, might be lost because it only relies on a single semantic feature in image reconstruction. Therefore, it is necessary to extract and transmit multi-modal semantic features from the original image to realize image reconstruction with high semantic similarity and satisfy the requirements of the receiver.
1.3 Image Generation-based Transmission Method with Multi-Modal Semantic Information
@@
To address this issue, we previously proposed a concept of an image transmission method that employs multiple types of semantic information [25]. In this method, a transmitter extracts multiple semantic features from an original image, such as composition and color, in addition to captions, and transmits them to a receiver, which reconstructs the image using an image-generation model. The images generated using multiple semantic features are expected to have a higher similarity to the original images than those generated using only a single feature. In this method, the receiver needs to evaluate the generated images based on some metrics and select an output result among the images. However, the specific procedures for the evaluation process and investigating evaluation metrics was not included in [25].
@@
In this paper, we propose a highly efficient image transmission method based on the concept proposed in [25] and employ newly designed evaluation metrics to assess the semantic similarity between the original and generated images. The proposed method aims to achieve monitoring, detecting and tracking specific objects mentioned in Sec. 1.2 in environments in which general image transmission is difficult described in Sec. 1.1. Fig. 1 shows an overview of the proposed method. In the proposed method, to achieve image transmission with a small amount of transmitted data, the transmitter extracts and transmits multi-modal semantic features of the original image, and the receiver reconstructs an image with the same semantic information as the received features. In the image generation process at the receiver, as the image-generation model generates various images, the receiver must select the best image as an output image. To evaluate the reconstruction result, this study developed quality assessment procedures, which is introduced later. In the proposed method, the receiver reconstructs images by the received semantic information to be easily recognized by humans, and therefore the proposed method can be potentially used in various applications.
@@
To select an appropriate image as an output in the proposed method, the quality of the generated images must be assessed based on the semantic similarity between the original and generated images on the receiver. Although some metrics assess the image quality by comparing two images [26], the proposed method cannot compare the original and generated images directly because the receiver does not possess the original image. Therefore, the quality of the generated images must be assessed based only on the received semantic features of the original image. To address this issue, this study proposes two scoring procedures for evaluating the similarity between semantic features, such as descriptive and segment information, of the original and generated images. Additionally, this study also proposes a background recoloring method for enhancing the contours of objects in images to improve image generation.
@@
  • We propose an image construction method that reconstructs images at the receiver end using multi-modal semantic communication and similarity evaluations. Therefore, the proposed method can transmit images with higher semantic similarity information than conventional transmission methods. Additionally, in this paper, we employed only open-source machine learning algorithms [27, 28, 20] for realizing the processes of the extraction of semantic features and image generation in the proposed method. This implementation highlights to facilitate the reproduction and adoption of the proposed method by others.
  • Extending from [25], this paper adds specific procedures to evaluate generated images by a receiver to select an output result among the images. The proposed method can select an image with high reproducibility as the output because the receiver evaluates the semantic similarity between the original and generated images to select the result. To realize this, we propose two scoring procedures for evaluating the semantic similarity between images. These procedures can evaluate the semantic similarity between images based only on the semantic features received by the receiver.
  • We propose and investigate scoring procedures of the semantic similarity between the original and generated images. This is an important first step toward establishing evaluation metrics for image transmission methods using semantic communication. The results of the experiment indicate that the proposed procedures can evaluate semantic features of the target object, such as position and composition, and background information of the image. Additionally, the results demonstrate that the proposed method with the background recoloring technique can effectively reconstruct the composition and position of the target object in the generated images.
@@ 2 Related Work
[24] proposed an image transmission method that transmits only the image caption to further improve communication efficiency. In this method, a transmitter generates a caption as descriptive information of the original image using an image captioning algorithm. Subsequently, the transmitter transmits only the generated caption, and the receiver reconstructs the image based on the received caption by using an image-generation model. As this method extracts only the descriptive information from the original image, it may incur a loss of some semantic information contained in the original image that is difficult to represent through the descriptive information. For example, it is difficult to represent the exact position and composition of a target object in an image using only descriptive information. Therefore, to reconstruct an image more clearly at the receiver, the transmitter must extract and transmit multi-modal semantic information from the original image.
To address this issue, we previously proposed a concept of novel image transmission method that uses multi-modal semantic information [25]. In this method, a transmitter extracts descriptive, segmented, and color information from the original image as semantic features, and transmits only these features. After receiving these features, the receiver reconstructs the image using an image-generation model. By conveying diverse information, such as the composition and color of the original image, in addition to descriptive information, this method enables image transmission that can generate images that are more semantically similar to the original image compared with a method using only single feature [24]. [25] included only concept of the proposed method and subjective classification of the generated images as a preliminary study. In the proposed method, the receiver needs to evaluate the generated images based on some metrics to select an output result among the images. However, specific procedures and investigating evaluation metrics for these evaluation and result selection processes are not included in [25].
@@ 3 Image Generation-based Transmission Method
This section describes the proposed image generation-based transmission method that uses multi-modal semantic information. A transmitter extracts multiple semantic features, such as composition, description, and color information, representing diverse semantic information from an image and transmits only these extracted features. The details of these extracted features are explained in Sec. 3.1. The receiver generates multiple images using the received features and selects an output image based on the semantic similarity. The details of these procedures at the receiver are explained in Sec. 3.2. The proposed method focuses on both a reduction of transmitted data size and image reconstruction while maintaining specific semantic features required by receiver’s application.
@@ 4 Evaluation Metrics for the Proposed Method
Table 1: Data sizes employed by each method. The average data size of the captions is approximately $1/270$ of that of the images in the JPEG format. Additionally, even when multiple semantic features are combined, the data size is approximately $1/14$ of that of the JPEG format images.
@@ 5 Experiment Setup
Refer to caption (a) Using only the caption. Refer to caption (b) Proposed method w/o background recoloring. Refer to caption (c) Proposed method with background recoloring.
@@
Refer to caption (a) Using only captions. Refer to caption (b) Proposed method w/o background recoloring. Refer to caption (c) Proposed method with background recoloring.
@@ 6.1 Data Sizes of Semantic Features
We compared the data size transmitted by the proposed method with that of the raw bit maps, JPEG formatted data, and generated captions of the original images to validate its communication efficiency. The reason we focused on transmission data size is that the proposed transmission method aims to realize a reduction of transmitted data size, and the reduction in transmission data size will enhance network performances such as shorter transmission delay and reduced network congestion. Therefore, this evaluation result indirectly confirms that effectiveness of the proposed method in environments where network resources are limited.
Table 1 presents the data sizes for each compression approach. “Uncompressed” denotes the size of the uncompressed bit map of the original images, “JPEG” denotes the average size of the original images in the JPEG format, and “Caption” denotes average size of the captions for each original image, which is used by the simple method employing only the descriptive information of the images. In addition, “Caption + color palette + segmentation array” denotes the average total size of the caption, color palette, and segmentation array for each image, which is the size of semantic features used in the proposed method. The sizes of all elements of the segmentation arrays and captions characters were calculated as 1 byte each. The data size of each color palette was calculated as the number of labels in an image $\times$ 4 bytes: 3 bytes for the RGB values and 1 byte for the label index. Each segmented array was compressed using run-length encoding for further data reduction. Additionally, “Caption + color palette + segmentation array” includes the RGB value of the “background” label.
The results indicate that the average data size of the captions is approximately $1/270$ of that of the images in the JPEG format. Additionally, even when multiple semantic features are combined, the data size is approximately $1/14$ of that of the JPEG format images. This result indicates that the proposed method can achieve higher efficiency in image transmission compared to general image compression algorithms. For example, LoRaWAN that is one of LPWA technologies provides data rate of up to 37.5 kbps in LoRa modulation [3]. This means that it takes approximately 8.9 seconds to transmit one JPEG image at this data rate. In contrast, it takes about 0.62 seconds to transmit semantic features equivalent to one image with the proposed method, thereby underscoring the utility of the proposed method in situations with limited network resources. Additionally, data size comparisons of the semantic features showed that only the caption extracted from the images had a smaller data size than the total size of the multiple semantic features, indicating that the communication efficiency was better when only captions were used. However, the proposed method was superior in terms of color and composition reproduction, as described in the following sections.
@@ 6.3 Comparison of Segmentation Matching Rate of Each Method
Fig. 8 shows the segmentation matching rates for the segmentation arrays generated from the original five “airplane” images and the 50 images generated using each method. Approximately, the same trends can be observed in the results of the other images, wherein the proposed methods exhibit higher matching rates than using only captions for the image generation. This is because the proposed methods can generate images with similar compositions and positions of the target object as those in the original image using the colored-segmented images.
@@ 7 Conclusion
This paper proposed an image-generation-based transmission method. In the proposed method, a transmitter extracts multi-modal semantic features from an image and transmits only them to a receiver. The receiver generates multiple images using an image-generation model and selects an output image based on the semantic similarity between the original and generated images. Therefore, the proposed method can realize both a significant reduction in the transmitted data size and reconstruction of the image with the semantic information required by the receiver. The evaluation results indicates that the proposed method can reduce the data size compared with typical image compression algorithms.
@@
In the proposed method, the receiver must evaluate the similarity between the original and generated images using only the received data to select the output. Therefore, this study proposed two scoring methods for comparing them between the original and generated images, and an experiment was conducted to verify the effectiveness of the proposed transmission method and scoring procedures. The results indicated that the proposed method using multi-modal semantic features can generate images that are more similar to the original images than those using only a single feature. In addition, the proposed scoring procedures can quantify the semantic similarity between images. The text similarity scores can compare the types of objects and rough background information in images and the segmentation matching rates can compare the composition and position of objects in the images. However, we found that if the target object in the image was too small, the generated images with no objects exhibited high segmentation matching rates.
@@
In a future work, we will improve the scoring procedures used for evaluating the semantic similarity. In the proposed scoring procedure for the segmentation matching rate, generated images with no objects had a high score when the target object in the original image was too small. This is because the matching rate is calculates based on the number of pixels including those labeled as “background.” Therefore, it is necessary to develop a calculation method that considers only the pixels labeled as “target object.” Moreover, we plan to develop an image reconstruction method based on the evaluation results of the semantic similarity. In the proposed method, the receiver selects the output from the generated images based on semantic similarity. However, if all generated images have low similarity scores, the receiver cannot select the result required by its application. To address this issue, it is necessary to develop a method to regenerate images or, if necessary, request the sender to transmit the semantic features again. Finally, we plan to apply the proposed transmission method to video streaming. The current proposed method does not focus on a processing delay caused by extraction of semantic features and image generation. However, this issue may be addressed by employing faster machine learning algorithms for real-time applications [23, 40, 41]. If the proposed transmission method can be applied to video streaming applications, this method will achieve a higher compression rate than conventional video compression algorithms.